OpenStudy (anonymous):

Does anyone know how to do imaginary numbers? I have one problem. that's a doozy

OpenStudy (anonymous):

doozy ? seriously . just give it to me im a genious! ok :)

OpenStudy (anonymous):

post it

OpenStudy (anonymous):

u know i aint got all day

OpenStudy (anonymous):

i kr

OpenStudy (anonymous):

HURRY!!!

OpenStudy (anonymous):

i have university soon

OpenStudy (anonymous):

ya know its advance

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

\[ \sqrt{3} \times \sqrt{-128}\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

guess ill be leaving if u dont give me the genius question some people are snobs

OpenStudy (anonymous):

i didnt get any younger u know

OpenStudy (anonymous):

please show steps. I have came up with three different answers

OpenStudy (anonymous):

post them

OpenStudy (anonymous):

@Tia1977 give me a medal first

OpenStudy (anonymous):

me 2

OpenStudy (anonymous):

then solve it as piece of cake

OpenStudy (anonymous):

would u like some cake it comes with genius ya know

OpenStudy (anonymous):

ok the answer is 48 U WELCOME!!1

OpenStudy (anonymous):

Thats impossible.<-- my answer

OpenStudy (anonymous):

well thx forr the medal, guess ill be leaving now, college ya know

OpenStudy (anonymous):

steps please!

OpenStudy (anonymous):

its 48

OpenStudy (anonymous):

first off, \[128=2^7\]

OpenStudy (anonymous):

sorry i dont do steps i do it all in myhead thats y im a genius

OpenStudy (anonymous):

see ya

OpenStudy (anonymous):

woudnt wanna be ya

OpenStudy (anonymous):

hahahaha that rhymes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and therefore \[\sqrt{-128}=\sqrt{2^7\times (-1)}=\sqrt{2^7}\times \sqrt{-1}=\sqrt{2^6}\times \sqrt{2}\times \sqrt{-1}=8\sqrt{2}i\]

OpenStudy (anonymous):

therefore \[\sqrt{3}\times 8\sqrt{2}i=8\sqrt{6}i\]

OpenStudy (anonymous):

ok. I got

OpenStudy (anonymous):

\[\sqrt{3} \pm \sqrt{2}\]

OpenStudy (anonymous):

not sure what is wrong